if I have a 32 in my English class and essays are worth 60 percent of my grade if I make a 100 on my essay what will be my new grade? what if I make a 90?80?70?

Answer by ewatrrr(24785)

.40*32 + .60*100 = 72.8
.40*32 + .60*90 = 66.8
etc
.40*32 + .60*80 = 48 + 12.8
.40*32 + .60*70 = 42 + 12.8